twice john's age, added to sylvia's, gives a total of 24.If the sum of John's age and Slyvia's age is 5 times john's age, how old is each?

Answer by ankor@dixie-net.com(22740)

twice john's age, added to Sylvia's, gives a total of 24.
2j + s = 24
:
If the sum of John's age and Sylvia's age is 5 times john's age,
j + s = 5j
subtract j from both sides
s = 5j - j
s = 4j
:
how old is each?
In the 1st equation, replace s with 4j
2j + 4j = 24
6j = 24
j = 24/6
j = 4 yrs is John's age
:
I'll let you find S's age, check your solutions in the 1st equation